Shift report 20011002 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Oct 02 2001 - 15:48:17 EDT

  • Next message: J.H. Lee: "dN/deta, Asymmetry, Background (Re: dndeta figures for impacients )"

    ------------------------------------------------------------
    Date: Tue Oct  2 15:48:17 2001, Shift: 08:00-16:00
    Supervisor: Eun-Joo Kim, with: 
    Log book#9 pages 75 - 76
    
    Shift Summary:
    ------------------------------------------------------------
    1. No beam : quench interlock, power dip etc
       "preparing to refill" message at 15:00
    
    2. No Achievements : pedestal, DVM run
    
    3. No detector problems.
    
       FV installed modified trigger/clock module.
       Inspected TPC pedestal run (run 5092), and
       nothing new from before.   
        
         
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Tue Oct 02 2001 - 15:48:46 EDT